Magic Realm, Inc., has developed a new fantasy board game. The company sold 27,900 games last year at a selling price of $65 per game. Fixed expenses associated with the game total $465,000 per year, and variable expenses are $45 per game. Production of the game is entrusted to a printing contractor. Variable expenses consist mostly of payments to this contractor.

Required:

1-a. Prepare a contribution format income statement for the game last year.

1-b. Compute the degree of operating leverage.

2. Management is confident that the company can sell 35,991 games next year (an increase of 8,091 games, or 29%, over last year). Given this assumption:

a. What is the expected percentage increase in net operating income for next year?

b. What is the expected amount of net operating income for next year? (Do not prepare an income statement; use the degree of operating leverage to compute your answer.)
Answer #1

1-a

Sales Revenue $          18,13,500.00 =27900*65
Less Variable Expenses $          12,55,500.00 =27900*45
Contribution Margin $             5,58,000.00
Less Fixed Expenses $             4,65,000.00
Net Operating Income $                93,000.00

1-b
Degree of Operating Leverage = Contribution Margin / Net Operating income
= $558000 / $93000 = 6

2-a
Expected Percentage increase in Net Operating Income = 29% x 6 = 174%

2-b
Expect amount of net operating income = $93000 + 174% = $254820